# Weakly Supervised Learning for Facial Behavior Analysis: A Review

### v3

@@ 6 Conclusion
This paper introduces various WSL scenarios for facial behavior analysis. A comprehensive review of state-of-the-art WSL methods for the recognition of expressions and AUs is provided for both classification and regression applications based on images and videos. In addition, the comparative results obtained with well-established experimental methodologies allow us to provide further insights into the benefits and limitations of these methods. Our review indicates that WSL represents a promising approach for training robust DL models for facial behavior analysis in real-world scenarios. Most of the WSL methods rely on either facial landmarks or pooling to capture the temporal dynamics, and efficient modeling of temporal dynamics with weak labels is expected to greatly improve system performance. Methods for deep WSL are still under-explored in the literature, especially for dimensional expression recognition, providing many opportunities for improving performance in WSL scenarios. Finally, we highlight the research gaps in the current literature pertinent to WSL models and provide potential research directions to advance the development of a robust system for facial behavior analysis with weak annotations.

### v4

@@ VI Conclusion
This paper introduces different WSL scenarios for facial behavior analysis. A comprehensive review of state-of-the-art WSL methods for the recognition of expressions and AUs is provided for both classification and regression applications based on images and videos. In addition, the comparative results obtained with well-established experimental methodologies allow us to provide further insights into the benefits and limitations of these methods. Our review highlights the diversity of WSL methods available for training robust DL models for facial behavior analysis in real-world scenarios. Most of the WSL methods rely on either facial landmarks or pooling to capture the temporal dynamics, and efficient modeling of temporal dynamics with weak labels is expected to greatly improve system performance. Methods for deep WSL are still under-explored in the literature, especially for dimensional expression recognition, providing many opportunities for improving performance in WSL scenarios. Finally, we highlight the research gaps in the literature on WSL models and provide potential research directions for the development of a robust system for facial behavior analysis with weak annotations.